Gdub08:

y=3(x-½)^2 +¼ the Vortex: Axis of Symmetry: Up or Down: Max or Min (value)

Gdub08:

Help me oml

Gdub08:

I need help @sailor

Glorious:

Vertex: (1/2, 1/4) Axis of Symmetry: x = 1/2 Up or Down: Up Max or Min: Min value is y = 1/4